# Copositive criteria for a two-component dark matter model

### v1

@@ 4 Dark matter phenomenology
$\mathcal{C}_{h_{1}S_{1}S_{1}}=-2i\cos\theta\lambda_{14}v_{0}+2i\sin\theta\lambda_{13}v_{1},$
$\mathcal{C}_{h_{2}S_{1}S_{1}}=-2i\sin\theta\lambda_{14}v_{0}-2i\cos\theta\lambda_{13}v_{1},$
$\mathcal{C}_{h_{1}S_{2}S_{2}}=-2i\cos\theta\lambda_{24}v_{0}+2i\cos\theta\lambda_{23}v_{1},$
$\mathcal{C}_{h_{2}S_{2}S_{2}}=-2i\sin\theta\lambda_{24}v_{0}-2i\cos\theta\lambda_{23}v_{1}$ (18)
Furthermore, the effective Lagrangian related to DM-nucleon elastic scattering can be given by:
$\mathcal{L}_{N,eff}=\sum_{S=S_{1},S_{2}}=\frac{m_{N}-\frac{7}{9}m_{B}}{v_{0}}(\frac{\mathcal{C}_{h_{1}SS}}{m_{h_{1}}^{2}}+\frac{\mathcal{C}_{h_{2}SS}}{m_{h_{2}}^{2}})SS\bar{N}N$ (19)
@@
where $m_{N}$ represents the nucleon mass and $m_{B}$ is the baryon mass in the chiral limit He:2008qm . The expression of the cross-section for the spin-independent DM- nucleon elastic scattering can thus be given as follows:
$\sigma_{S_{i}N}=\frac{m_{N}^{4}f_{N}^{2}}{4\pi(m_{N}+m_{i})^{2}}\times(\frac{\mathcal{C}_{h_{1}S_{i}S_{i}}}{m_{h_{1}}^{2}}+\frac{\mathcal{C}_{h_{2}S_{i}S_{i}}}{m_{h_{2}}^{2}})^{2},(i=1,2)$ (20)
@@
where $f_{N}$ is the Higgs-nucleon Form factor with $f_{N}=0.308(18)$ according to the phenomenological and lattice-QCD calculations Bell:2013wua . Since we have two dark matter particles in the model, we should compare the value of $\xi_{i}\sigma_{S_{i}N}$ against the direct detection constraints provided by the experiment results instead of the cross-section, where $\xi_{i}$ is the fraction of $S_{i}$ in the total relic density defined by:
$\xi_{i}=\frac{\Omega_{i}}{\Omega_{DM}},(i=1,2)$ (21)
@@
and $\Omega_{DM}$ is the observed value of the dark matter density $\Omega_{i}$ is the density of $S_{i}(i=1,2)$. In our analysis, we will consider the current direct detection limit set by the XENONnT XENON:2015gkh and LZ LZ:2022lsv .
5 Discussion
@@
In this part, we estimate the parameter space from the view of copositive criteria and dark matter phenomenology related to relic density and direct detection constraints. As we mentioned above, there are eight cases related to copositive criteria that depend on the sign of the parameters, and each case corresponds to a viable parameter space that should be limited by the relic density constraint and direct detection constraint. On the other hand, regardless of the symmetric matrix $\mathcal{S}$ with the permutation $\lambda_{ij}=\lambda_{ji}$, we can still have different results when the indexes $i,j,k,l$ takes different permutations even just for one case. Hence, discussion about the parameter space will be complex for different permutations of the indexes and one should consider all the possible cases to obtain the viable parameter space.
5.1 Possible permutations of the indexes
To decrease the complexity and simplify the discussion, we consider the following assumptions. Firstly, notice that we have an exchange symmetry of $\lambda_{13}\leftrightarrow\lambda_{23}$, $\lambda_{14}\leftrightarrow\lambda_{24}$ and $S_{1}\leftrightarrow S_{2}$ due to the $Z_{2}\times Z_{2}^{\prime}$ symmetry, and there is no preference for the permutations of the indexes when considering the copositive criteria. Therefore, we will only consider one possible case when discussing the related parameters since we will have a similar shape of the parameter space. On the other hand, $\lambda_{34}$ is the coupling of the new Higgs particles, which makes little difference in dark matter results, and we omit it as a free parameter and will not consider the cases including the sign of $\lambda_{34}$ in the following discussion.
Taking the above assumptions into consideration, we can conclude all the possible cases as follows when discussing copositive criteria:
I) For Case I with $\lambda_{ij}\geqslant 0$, we have all the couplings larger than 0.
II) For case II with $\lambda_{ij}\leqslant 0$, we have the following permutations: $\lambda_{12}\leqslant 0$ or $\lambda_{13}\leqslant 0$ or $\lambda_{24}\leqslant 0$ or $\lambda_{14}\leqslant 0$ or $\lambda_{23}\leqslant 0$. As we mentioned above, we ignore the case $\lambda_{34}\leqslant 0$. Meanwhile, for the cases $\lambda_{13}\leqslant 0$ and $\lambda_{23}\leqslant 0$, we only need to consider one of these two cases due to the exchange symmetry and here we can consider $\lambda_{13}\leqslant 0$ for the following discussion. Similarly, for $\lambda_{14}\leqslant 0$ and $\lambda_{24}\leqslant 0$ we consider $\lambda_{24}\leqslant 0$. Hence, there are 3 cases to be discussed for Case II with $\lambda_{12}\leqslant 0$ or $\lambda_{13}\leqslant 0$ or $\lambda_{14}\leqslant 0$.
III) For Case III with $\lambda_{ij},\lambda_{lk}\leqslant 0$, there are $C^{2}_{4}=3$ kinds of permutations in total, but we will only need to discuss the case with $\lambda_{14},\lambda_{23}\leqslant 0$ and the case $\lambda_{12},\lambda_{34}\leqslant 0$ is ignored while $\lambda_{13},\lambda_{24}\leqslant 0$ has the similar results with $\lambda_{14},\lambda_{23}\leqslant 0$.
IV) For Case IV with $\lambda_{ij},\lambda_{ik}\leqslant 0$, the amount of the possible permutations is $\frac{4\times 3\times 2}{2}-2\times 2=8$ without the cases including the sign of $\lambda_{34}$, and we will have 5 different kinds of permutations in total after taking exchange symmetry into consideration. Concretely speaking, we will discuss the cases with $\lambda_{12},\lambda_{13}\leqslant 0$,$\lambda_{12},\lambda_{14}\leqslant 0$,$\lambda_{13},\lambda_{14}\leqslant 0$,$\lambda_{23},\lambda_{13}\leqslant 0$,$\lambda_{14},\lambda_{24}\leqslant 0$
V) For Case V with $\lambda_{ij},\lambda_{jk},\lambda_{ik}\leqslant 0$, we have $\frac{4\times 3\times 2}{2\times 2\times 2}=3$ kinds of different permutations, and we have $\lambda_{12},\lambda_{23},\lambda_{13}\leqslant 0$ as well as $\lambda_{12},\lambda_{24},\lambda_{14}\leqslant 0$ after ignoring the case including $\lambda_{34}$.
VI) For Case VI with $\lambda_{ij},\lambda_{ik},\lambda_{il}\leqslant 0$ we have $\lambda_{12},\lambda_{13},\lambda_{14}\leqslant 0$ after taking exchange symmetry into consideration.
VII) For Case VII with $\lambda_{ij},\lambda_{jk},\lambda_{kl}\leqslant 0$, we have 3 different kinds of permutations left after taking exchange symmetry into consideration and ignoring the case including $\lambda_{34}$. Concretely speaking, we have $\lambda_{13},\lambda_{23},\lambda_{24}\leqslant 0$, $\lambda_{14},\lambda_{24},\lambda_{23}\leqslant 0$ and $\lambda_{12},\lambda_{13},\lambda_{24}\leqslant 0$
VIII) For Case VIII with $\lambda_{ij},\lambda_{jk},\lambda_{kl},\lambda_{il}\leqslant 0$, we have $\frac{4\times 3\times 2}{2\times 4}=3$ kinds of permutations, and we only need to discuss the case $\lambda_{13},\lambda_{23},\lambda_{24},\lambda_{14}\leqslant 0$ after considering exchange symmetry and ignoring the case including $\lambda_{34}$.
In conclusion, we will consider $17(1+3+1+5+2+1+3+1)$ different kinds of permutations in total in the following discussion.
5.2 Theortical constraint of copositive criteria
Following the above discussion, we consider the 17 different cases with theoretical constraint of copositive criteria on the viable parameter space. In this work, we focus on the parameters related to dark matter phenomenology and copositive criteria. We can choose the following parameters as inputs:
$v_{1},\sin\theta,m_{h_{2}},m_{1,2},\lambda_{11},\lambda_{12},\lambda_{13},\lambda_{14},\lambda_{22},\lambda_{23},\lambda_{24}$ (22)
while $\lambda_{33}$,$\lambda_{34}$ and $\lambda_{44}(\lambda_{H})$ can be given by Eq. 2. Note that for the large couplings, the parameter space can be excluded by direct detection constraints, and in the following discussion we assume the couplings to be smaller than $\pi$ for simplicity. On the other hand, since the couplings $\lambda_{ii}$ are less relevant with dark matter phenomenology with $i=1,2,3,4$ and one can always fine-tune these couplings to satisfy the related inequalities to guarantee the copositive criteria.
For Case I, Case II and Case III, the copositive criteria can be satisfied by fine-tuning the $\lambda_{ii}$ with $i=1,2,3,4$ as we mentioned above. For Case IV, the parameter space is constrained by the inequality to satisfy the copositive criteria. In Fig. 1(a), we give the result of Case IV, where the colored region is the viable parameter space of $\lambda_{ij}-\lambda_{ik}$ satisfying the copositive criteria and we have set $\lambda_{ii}=\lambda_{jj}=\lambda_{kk}=\pi$ and $\lambda_{jk}=1$. The small negative $\lambda_{ij}-\lambda_{ik}$ pair is excluded for the copositive criteria and for a smaller $\lambda_{jk}$ the allowed parameter space for $(\lambda_{ij},\lambda_{ik})$ is more stringent according to copositive criteria.
According to Case V, the copositive criteria demands the symmetric matrix of order three is copositive, and we have:
$\lambda_{ij}+\sqrt{\lambda_{ii}\lambda_{jj}}\geqslant 0,\lambda_{ik}+\sqrt{\lambda_{ii}\lambda_{kk}}\geqslant 0,\lambda_{jk}+\sqrt{\lambda_{jj}\lambda_{kk}}\geqslant 0,\sqrt{\lambda_{ii}\lambda_{jj}\lambda_{kk}}+\lambda_{ij}\sqrt{\lambda_{kk}}+\lambda_{ik}\sqrt{\lambda_{jj}}$
$+\lambda_{jk}\sqrt{\lambda_{ii}}+\sqrt{2(\lambda_{ij}+\sqrt{\lambda_{ii}\lambda_{jj}})(\lambda_{ik}+\sqrt{\lambda_{ii}\lambda_{kk}})(\lambda_{jk}+\sqrt{\lambda_{jj}\lambda_{kk}})}\geqslant 0$ (23)
Figure 1: The allowed parameter space constrained by the copositive criteria corresponds to the colored region, where (a) represents the result of Case IV, (b) corresponds to the result of Case V, (c) corresponds to Case VI, (d) is the result of Case VII and (e) corresponds to Case VIII.
We give the result of Case V in Fig. 1(b), and the colored region is the viable parameter space of $\lambda_{ij}-\lambda_{ik}$ satisfying the copositive criteria, where we have set $\lambda_{ii}=\lambda_{jj}=\lambda_{kk}=\pi$ and $\lambda_{jk}=-1$. For the negative $\lambda_{ij}-\lambda_{ik}$, the left-bottom region corresponding to small $\lambda_{ij}-\lambda_{ik}$ is excluded with the copositive criteria.

### v4

@@ IV Dark matter phenomenology
$\mathcal{C}_{h_{1}S_{1}S_{1}}=-2\cos\theta\lambda_{14}v_{0}+2\sin\theta\lambda_{13}v_{1},$
$\mathcal{C}_{h_{2}S_{1}S_{1}}=-2\sin\theta\lambda_{14}v_{0}-2\cos\theta\lambda_{13}v_{1},$
$\mathcal{C}_{h_{1}S_{2}S_{2}}=-2\cos\theta\lambda_{24}v_{0}+2\sin\theta\lambda_{23}v_{1},$
$\mathcal{C}_{h_{2}S_{2}S_{2}}=-2\sin\theta\lambda_{24}v_{0}-2\cos\theta\lambda_{23}v_{1}.$ (18)
@@
The expression of the cross-section for the spin-independent DM- nucleon elastic scattering can thus be given as follows He et al. (2009):
$\sigma_{S_{i}N}=\frac{m_{N}^{4}f_{N}^{2}}{4\pi(m_{N}+m_{i})^{2}}\times(\frac{\mathcal{C}_{h_{1}S_{i}S_{i}}}{m_{h_{1}}^{2}}\cos\theta+\frac{\mathcal{C}_{h_{2}S_{i}S_{i}}}{m_{h_{2}}^{2}}\sin\theta)^{2},(i=1,2)$ (19)
@@
where $m_{N}$ represents the nucleon mass and $f_{N}$ is the Higgs-nucleon Form factor with $f_{N}=0.308(18)$ according to the phenomenological and lattice-QCD calculations Bell et al. (2014). Since we have two dark matter particles in the model, we should compare the value of $\xi_{i}\sigma_{S_{i}N}$ against the direct detection constraints provided by the experiment results instead of the cross-section, where $\xi_{i}$ is the fraction of $S_{i}$ in the total relic density defined by:
$\xi_{i}=\frac{\Omega_{i}}{\Omega_{DM}},(i=1,2).$ (20)
@@
and $\Omega_{DM}$ is the observed value of the dark matter density $\Omega_{i}$ is the density of $S_{i}(i=1,2)$. In our analysis, we will consider the current direct detection limit set by the LZ result Aalbers and others (2024).
V Discussion
@@
In this part, we estimate the parameter space from the view of copositive criteria and dark matter phenomenology related to relic density and direct detection constraints. Provided that direct detection constraint is more stringent on the low mass region, we focus on the TeV scale dark matter mass in this work, where indirect detection constraints make little difference on the parameter space. As we mentioned above, there are eight cases related to copositive criteria that depend on the sign of the parameters, and each case corresponds to a viable parameter space that should be limited by the relic density constraint and direct detection constraint. On the other hand, regardless of the symmetric matrix $\mathcal{S}$ with the permutation $\lambda_{ij}=\lambda_{ji}$, we can still have different results when the indexes $i,j,k,l$ take different permutations even just for one case. Hence, discussion about the parameter space will be complex for different permutations of the indexes and one should consider all the possible cases to obtain the viable parameter space.
V.1 Estimate on the parameters
Before discussion, we can estimate the parameters to simplify the discussion. In this work, we choose the following parameters as inputs:
$v_{1},\sin\theta,m_{h_{2}},m_{1,2},\lambda_{11},\lambda_{12},\lambda_{13},\lambda_{14},\lambda_{22},\lambda_{23},\lambda_{24}.$ (21)
while $\lambda_{33}$$\lambda_{34}$ and $\lambda_{44}(\lambda_{H})$ can be given by Eq. II. Note that $\lambda_{34}$ is the coupling of the new Higgs particles, which makes little difference in dark matter results, and one can fix $v_{1}$, $\sin\theta$ as well as $m_{h_{2}}$ so that $\lambda_{33},\lambda_{34}$ and $\lambda_{44}(\lambda_{H})$ can be given explicitly. Note that the self-interaction couplings $\lambda_{11}$ and $\lambda_{22}$ make little difference in dark matter production, and we fix $\lambda_{11}=\lambda_{22}=\pi$ for simplicity. Therefore, we can just consider the signs of $\lambda_{12},\lambda_{13},\lambda_{23},\lambda_{14},\lambda_{24}$. Notice that we have an exchange symmetry of $\lambda_{13}\leftrightarrow\lambda_{23}$, $\lambda_{14}\leftrightarrow\lambda_{24}$ and $S_{1}\leftrightarrow S_{2}$ due to the $Z_{2}\times Z_{2}^{\prime}\times Z_{2}^{\prime\prime}$ symmetry and there is no preference for the permutations of the indexes, one can group the five couplings into $\lambda_{12}$$(\lambda_{13},\lambda_{14})$ and $(\lambda_{23},\lambda_{24})$.
[…]
Figure 2: Viable parameter space of $(\lambda_{13},\lambda_{14})$ under direct detection constraint in the absence of $S_{2}$ with $m_{1}=1$ TeV.
V.2 Theortical constraint of copositive criteria
Following the above discussion, we have 4 different cases left with the theoretical constraints of copositive criteria to discuss. The case (1) corresponds to Case I in Sec III where all the couplings are larger than zero. The case (2) corresponds to Case IV with $\lambda_{13},\lambda_{14}\leqslant 0$, and the copositive criterial demands that:
$\lambda_{11}\lambda_{34}-\lambda_{13}\lambda_{14}+\sqrt{(\lambda_{11}\lambda_{33}-\lambda_{13}^{2})(\lambda_{11}\lambda_{44}-\lambda_{14}^{2})}\geqslant 0,$ (22)